Two functions f : R → R and g : R → R are defined by f(x) = 5x3 + 1 and g(x) = 2x − 3 for all x ∈ R.

Determine the inverse of (f -1 ◦ g) and (g ◦ f )(2) and ( f ◦ g)(2).


1
Expert's answer
2022-05-26T08:04:15-0400

Let "f(x_1)=f(x_2)." It means that



"5x_1^3+1=5x_2^3+1""x_1^3=x_2^3""(x_1-x_2)(x_1^2+x_1x_2+x_3^2)=0""x_1-x_2=0""x_1=x_2"

The function "f(x)=5x^3+1" is bijective (one-to-one ) from "\R" to "\R."



"f(x)=5x^3+1, x\in \R""y=5x^3+1"

Change "x" and "y"

"x=5y^3+1"

Solve for "y"

"y^3=\dfrac{x-1}{5}"




"y=\sqrt[3]{\dfrac{x-1}{5}}"

Then


"f^{-1}(x)=\sqrt[3]{\dfrac{x-1}{5}}, x\in \R"



a)


"(f^{-1}\circ g)=\sqrt[3]{\dfrac{2x-3-1}{5}}=\sqrt[3]{\dfrac{2x-4}{5}}, x\in \R"


"y=\sqrt[3]{\dfrac{2x-4}{5}}"

Change "x" and "y"

"x=\sqrt[3]{\dfrac{2y-4}{5}}"

Solve for "y"

"2y-4=5x^3"




"y=\dfrac{5}{2}x^3+2"

Then


"(f^{-1}\circ g)^{-1}=\dfrac{5}{2}x^3+2, x\in \R"



b)


"(g\circ f)(x)=2(5x^3+1)-3=10x^3-1"




"(g\circ f)(2)=10(2)^3-1=79"

c)


"(f\circ g)(x)=5(2x-3)^3+1"

"(f\circ g)(2)=5(2(2)-3)^3+1=6"